OpenStudy (anonymous):

Write the product in standard form. (2-3i)(2+i)

OpenStudy (maheshmeghwal9):

\[(4+2i-6i-3i^2)=(4-4i+3)=(7-4i)\]

OpenStudy (maheshmeghwal9):

only multiply each term of one complex no. with another's:)